4) You have found the following information on a stock option: Stock price = $60; strike price = $65, Call price = $3. The option expires in 6 months, and the current risk-free rate is 3.1%. Calculate the option put price?
Answer #1

Using Put Call Parity Equation,

C + X/(1 + r)n = Stock Price + P

C = Price of Call Option

X = Exercise Price

P = Price of Put Option

3 + 65/(1.031)6/12 = 60 + P

P = $7.02

So,

Price of Put Option = $7.02